OpenStudy (anonymous):

what is the value of x?? (2x+14)(x+14)=952

OpenStudy (turingtest):

FOIL: 2x^2+2(14x)+14x+196=2x^2+42x+196=952 2x^2+42x-756=x^2+21x-378=0 can you use the quadratic formula from here? 'cause I don't want to...

OpenStudy (turingtest):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (unklerhaukus):

(2x-14)(x-14)=9522 2(x-7)(x-14)=9522 (x-7)(x-14)=4761

OpenStudy (turingtest):

where \[ax^2+bx+c=0\] is the form so plug in a=1 b=21 c=-378 into the formula above and simplify

OpenStudy (anonymous):

wow you guys are math wizs:)

OpenStudy (turingtest):

it's just algebra :)

OpenStudy (turingtest):

@ unk dude i think you messed up and forgot about the -4761, check your answer.

OpenStudy (turingtest):

a=1 b=21 c=-378 \[x=(-21 \pm \sqrt{21^2-4(1)(-378)})/2=(-21 \pm \sqrt{1953)}/2\]

OpenStudy (turingtest):

\[=(-21 \pm 3\sqrt{217})/2=(3/2)(-7 \pm \sqrt{217)}\]

OpenStudy (turingtest):

so\[x=\left\{ (3/2)(-7-\sqrt{217},(3/2)(-7+\sqrt{217)} \right\}\] is what I got

OpenStudy (unklerhaukus):

whoops

OpenStudy (turingtest):

am I right? or did I mess up too? this one was long and annoying :(

OpenStudy (unklerhaukus):

im gonna try on on paper instead of in my head...

OpenStudy (turingtest):

@ alexandraox3 I think my answer is the right one.